# The effect of priors on Learning with Restricted Boltzmann Machines

### v1

@@ Conclusions
We show that continuous (Gaussian) hidden units help the machine to more easily enter the signal retrieval region, even though the critical size only depends on the properties of the dataset, i.e., the teacher RBM. At the same time, certain choices of priors favor the emergence of an example retrieval region where the machine learns by memorization without generalizing well. Notably, the posterior distribution approach for training includes automatic self-regularization of the weights, preventing parameter explosions and other typical convergence issues.
Acknowledgements
@@ Appendix A Derivation of the RS equations
$Z^{n}(\bm{\mathcal{S}})=\mathbb{E}_{\{\bm{\xi}^{\gamma}\}}z(\{\bm{\xi}^{\gamma}\})^{-M}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\exp\left(M\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}p_{\tau}^{\gamma}(\tau)\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}p^{\gamma}(\xi)\tau^{\gamma}_{(1)}+\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}\sum_{i=2}^{N}\sum_{a=2}^{M}{s}_{i}^{a}\tau^{a,\gamma}\xi_{i}^{\gamma}\right)\,.$ (61)
The shorthand notation $\{\bm{\xi}^{\gamma}\}$ stands for the average over all the replicas indexed by $\gamma$: $\{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}\}$, same is done for $\{\bm{\tau}^{\gamma}\}$ Then the averaged partition function over the quenched disorder induced by the dataset is
$\left[Z^{n}\right]^{\bm{\mathcal{S}}}=\sum_{\bm{\mathcal{S}}}P_{\hat{\beta}}(\bm{\mathcal{S}})\ Z^{n}(\bm{\mathcal{S}})$ $=\sum_{\bm{\mathcal{S}}}\,\mathbb{E}_{\hat{\bm{\xi}}}\mathbb{E}_{\hat{\bm{\tau}}}\ z(\hat{\bm{\xi}})^{-M}\ e^{\sqrt{\hat{\beta}/N}\sum_{ia}s_{i}^{a}\hat{\xi}_{i}\hat{\tau}_{a}}\ Z^{n}(\bm{\mathcal{S}})\,,$ (62)
@@
The presence of both $z(\hat{\bm{\xi}})$ and $z(\bm{\xi})$ make it diffucukt to compute $\left[Z^{n}\right]^{\bm{\mathcal{S}}}$ as a straightforward extremization problem. A possible solution is the application of the saddle point method to both (63) and (58). We can write them as an exponential function to be extremized over a set of parameters $\bm{p}_{T}$ and $\bm{p}_{S}$. Each set will depend respectively on $\bm{\hat{\xi}}$ and $\bm{\xi}$
$z(\bm{\hat{\xi}})^{M}\approx e^{-N\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}\hat{\beta}f_{T}(\bm{p}_{T}(\bm{\hat{\xi}}))}\,,\qquad\qquad z(\bm{\xi})^{M}\approx e^{-N\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}\beta f_{S}(\bm{p}_{S}(\bm{\xi}))}\,,$ (64)
@@
$z(\hat{\bm{\xi}})^{M}$ $=$ $\int\prod_{a}\left[\mathcal{D}{\bm{g}}^{a}\right]\,\mathbb{E}_{\bm{\hat{\tau}}}\exp\left\{\sqrt{\frac{\hat{\beta}}{N}}\sum_{i=1}^{N}\sum_{a=1}^{M}\sqrt{\Omega_{s}}\,g_{s,i}^{a}\hat{\tau}^{a}\hat{\xi}_{i}\right\}\prod_{i,a}2\cosh\left\{\sqrt{\frac{\hat{\beta}}{N}}\sqrt{\delta_{s}}\hat{\tau}^{a}\hat{\xi}_{i}\right\}$ (65)
$\approx$ $\mathbb{E}_{\bm{\hat{\tau}}}\,\exp\left\{\frac{\hat{\beta}}{2N}\sum_{i=1}^{N}\sum_{a=1}^{M}\left(\hat{\tau}_{a}\hat{\xi}_{i}\right)^{2}\right\}\,,$ (66)
@@
[…]
@@
[…]
$\approx$ $e^{-N\hat{\beta}\operatorname{Extr}_{\bm{p}_{T}}f(\bm{p}_{T})}\,,$ (70)
@@
$-\hat{\beta}f(\bm{p}_{T})=\alpha\left(-d_{\hat{\tau}}\hat{d}_{\hat{\tau}}+\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\hat{d}_{\hat{\tau}}+\log\mathbb{E}_{\hat{\tau}}e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\right)$ (71)
@@
$\hat{d}_{\hat{\tau}}=\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\,,\qquad\qquad d_{\hat{\tau}}=\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}=\frac{\mathbb{E}_{\hat{\tau}}\hat{\tau}^{2}e^{\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\hat{\tau}^{2}}}{\mathbb{E}_{\hat{\tau}}e^{\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\hat{\tau}^{2}}}\,.$ (72)
@@
$\left[Z^{n}\right]^{\bm{\mathcal{S}}}$ $=$ $\sum_{\bm{\mathcal{S}}}\,\mathbb{E}_{\hat{\bm{\xi}}}\mathbb{E}_{\hat{\bm{\tau}}}\ e^{N\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f(\bm{p}_{T})}\ e^{\sqrt{\hat{\beta}/N}\sum_{ia}s_{i}^{a}\hat{\xi}_{i}\hat{\tau}_{a}}\ Z^{n}(\bm{\mathcal{S}})$ (73)
@@
The same problem of the teacher normalization term affects Eq.(61), due to the presence of $z(\{\bm{\xi}^{\gamma}\})$. As already stated in Eqs.(64) we want to write it in a variational principle manner. The steps to do that are the same of the one used for (65-72). We start computing the sum over the examples using their interpolated form, then we impose the self overlap with the Dirac delta
$z(\{\bm{\xi}^{\gamma}\})^{M}$ $=$ $\int\prod_{a,\gamma}\left[\mathcal{D}\bm{g}^{a\gamma}\right]\,\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\exp\left\{\sqrt{\frac{\beta}{N}}\sum_{\gamma,i,a}\sqrt{\Omega_{s}}\,g_{s,i}^{a\gamma}{\tau}_{a}^{\gamma}\xi_{i}^{\gamma}\right\}\prod_{i,a,\gamma}2\cosh\left\{\sqrt{\frac{\beta}{N}}\sqrt{\delta_{s}}\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right\}$ (74)
$\approx$ $\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\,\exp\left\{\frac{\beta}{2N}\sum_{\gamma=1}^{n}\sum_{i=1}^{N}\sum_{a=1}^{M}\left(\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right)^{2}\right\}\,,$ (75)
$\approx$ $\int\prod_{\gamma}\,dd_{\tau}^{\gamma}\,\mathcal{D}(\bm{\hat{d}_{\tau}},\bm{d_{\tau}})\exp\left\{\frac{\alpha\beta}{2}\sum_{\gamma}d_{\tau}^{\gamma}d^{\gamma}({\bm{\xi}})\right\}$ (76)
$\approx$ $e^{-N\beta\operatorname{Extr}_{\bm{p}_{S}}f(\bm{p}_{S})}\,,$ (77)
@@
$\mathcal{D}(\bm{\hat{d}_{\tau}},\bm{d_{\tau}})$ $=$ $\sum_{\{\bm{\tau}^{\gamma}\}}\delta\left(\sum_{\gamma}d_{{\tau}}^{\gamma}-\frac{1}{M}\sum_{a,\gamma}(\tau_{a}^{\gamma})^{2}\right)\propto\sum_{\{\bm{\tau}^{\gamma}\}}\int\prod_{\gamma}d\hat{d}_{\tau}^{\gamma}\,e^{-M\sum_{\gamma}\hat{d}_{\tau}^{\gamma}\left(d_{\tau}^{\gamma}-\frac{1}{M}\sum_{a}(\tau_{a}^{\gamma})^{2}\right)}\,,$
@@
$Z^{n}(\bm{\mathcal{S}})$ $=$ $\mathbb{E}_{\{\bm{\xi}^{\gamma}\}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\exp\left(M\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}p_{\tau}^{\gamma}(\tau)\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}p^{\gamma}(\xi)\tau^{\gamma}_{(1)}+\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}\sum_{i=2}^{N}\sum_{a=2}^{M}{s}_{i}^{a}\tau^{a,\gamma}\xi_{i}^{\gamma}+N\beta\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}f_{S}(\bm{p}_{S})\right)\,.$
@@
$-\beta f_{S}(\bm{p}_{S})$ $=$ $-\alpha\sum_{\gamma}d_{{\tau}}^{\gamma}\hat{d}_{{\tau}}^{\gamma}+\frac{\alpha\beta}{2}\sum_{\gamma}d^{\gamma}({\bm{\xi}}){d}_{{\tau}}^{\gamma}+\frac{1}{N}\log\mathbb{E}_{\{\tau^{\gamma}_{a}\}}e^{\sum_{\gamma,a}\hat{d}_{{\tau}}^{\gamma}\sum_{a}(\tau^{\gamma}_{a})^{2}}$ (78)
$\hat{d}_{{\tau}}^{\gamma}=\frac{\beta}{2}d^{\gamma}({\bm{\xi}})\,,\qquad d^{\gamma}_{\hat{\tau}}=\langle{(\tau^{\gamma})}^{2}\rangle_{\tau}=\frac{\mathbb{E}_{\tau^{\gamma}}(\tau^{\gamma})^{2}\,e^{\frac{\beta}{2}d^{\gamma}(\bm{\xi})(\tau^{\gamma})^{2}}}{\mathbb{E}_{\tau^{\gamma}}e^{\frac{\beta}{2}d^{\gamma}(\bm{\xi})(\tau^{\gamma})^{2}}}\,.$ (79)
RS free energy ans saddle point equations
@@
$\left[Z^{n}\right]^{\bm{\mathcal{S}}}=$$\sum_{\{\xi_{(1)}^{\gamma}\}}\mathbb{E}_{\{\tau_{(1)}^{\gamma}\}}\sum_{\bm{s},\,\tilde{\bm{s}}}\exp{\left\{M\sqrt{\frac{\beta}{N}}p_{\tau}^{\gamma}(\tau)\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}p(\xi)^{\gamma}\tau^{\gamma}_{(1)}\right\}}\times$ (80)
$\times$ $\mathbb{E}_{\hat{\bm{\xi}}}\mathbb{E}_{\hat{\bm{\tau}}}\sum_{\bm{\mathcal{S}}}\exp\left\{N\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f_{T}(\bm{p}_{T})+\sqrt{\frac{\hat{\beta}}{N}}\sum_{i=2}^{N}\sum_{a=2}^{M}{s}_{i}^{a}\hat{\tau}^{a}\hat{\xi}_{i}\right\}\times$ (81)
$\times$ $\sum_{\{\bm{\xi}^{\gamma}\}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\;\exp\left\{N\beta\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}f_{S}(\bm{p}_{S})+\sqrt{\frac{\beta}{N}}\sum_{i=2}^{N}\sum_{a=2}^{M}\sum_{\gamma=1}^{n}{s}_{i}^{a}\tau^{a,\gamma}\xi_{i}^{\gamma}\right\}\;.$ (82)
@@
$\int\,\prod_{\gamma}dp^{\gamma}\,\prod_{\gamma}dp_{\tau}^{\gamma}\,\mathcal{D}(\bm{p},\bm{p}_{\tau})\sum_{\{\xi_{(1)}^{\gamma}\}}\mathbb{E}_{\{\tau_{(1)}^{\gamma}\}}\exp{\left\{M\sqrt{\frac{\beta}{N}}\sum_{\gamma}p_{\tau}^{\gamma}\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}\sum_{\gamma}p^{\gamma}\tau^{\gamma}_{(1)}\right\}}$
$=$ $\int\,\prod_{\gamma}dp^{\gamma}\,\prod_{\gamma}dp_{\tau}^{\gamma}\,\mathcal{D}(\bm{p},\bm{p}_{\tau})\exp{\left\{\frac{N\beta}{2}\alpha^{2}\Omega\sum_{\gamma}(p_{\tau}^{\gamma})^{2}+\frac{N\beta}{2}\Omega_{\tau}\sum_{\gamma}(p^{\gamma})^{2}\right\}}\,,$ (83)
@@
$\propto\sum_{\bm{s},\,\tilde{\bm{s}}}\int\,\prod_{\gamma}d\hat{p}^{\gamma}\prod_{\gamma}d\hat{p}_{\tau}^{\gamma}\exp\left(-N\sum_{\gamma}\hat{p}_{\gamma}\left(p_{\gamma}-\frac{1}{N}\sum_{i}s_{i}\xi^{\gamma}_{i}\right)\right)-\left(M\sum_{\gamma}\hat{p}^{\gamma}_{\tau}\left(p^{\gamma}_{\tau}-\frac{1}{M}\sum_{a}\tilde{s}_{a}\tau^{\gamma}_{a}\right)\right)\;.$

### v2

@@ 6 Conclusions
We show that continuous (Gaussian) hidden units help the machine to more easily enter the signal retrieval region, even though the critical size only depends on the properties of the dataset, i.e., the teacher RBM. At the same time, certain choices of priors favor the emergence of an example retrieval region where the machine learns by memorization without generalizing well. Notably, the posterior distribution approach for training includes automatic self-regularization of the weights, preventing parameter explosions and other typical convergence issues. By combining replica analysis with numerical simulations, we explicitly tested how the choice of priors for the hidden units affects learning performance, confirming that Gaussian priors lead to better signal retrieval with a smaller data requirement. Although the teacher’s parameters are unknown in practice, the mismatched analysis provides useful insight into which student architectures—such as those using Gaussian units—tend to be more robust across different dataset properties. In this sense, our framework can guide the identification of architectural choices that generalize well and help establish lower bounds on dataset size required for successful learning, independently of the precise nature of the data-generating process.
A promising extension of this work would be to incorporate more structured data-generating processes, such as those involving correlated planted signals. Prior studies have explored such settings in specific regimes, e.g. RBM with binary synapses in hou2019minimal , theriault2024modelling , showing the occurrence of permutation symmetry breaking phase transitions. Similar transitions are observed in bachtis2024cascade , where the training dynamics on a dataset generated from correlated patterns has been analyzed. Recent approaches xu2025learning introduce a structure in the dataset by directly using a spiked covariance model as generating model for the examples. Extending our framework, which smoothly interpolates between Gaussian and binary priors, to these structured scenarios could reveal universal behaviors in feature-learning dynamics across the full class of RBMs. Such analysis would offer deeper insight into the interplay between data structure, noise, and architectural choices in shaping generalization performance in more realistic learning settings.
@@ Appendix A Derivation of the RS equations
$\hskip-28.45274ptZ^{n}(\bm{\mathcal{S}})=\mathbb{E}_{\{\bm{\xi}^{\gamma}\}}z(\{\bm{\xi}^{\gamma}\})^{-M}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\exp\Bigg($ $M\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}p_{\tau}^{\gamma}(\tau)\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}p^{\gamma}(\xi)\tau^{\gamma}_{(1)}$ (58)
$+\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}\sum_{i=2}^{N}\sum_{a=2}^{M}{s}_{i}^{a}\tau^{a,\gamma}\xi_{i}^{\gamma}\Bigg)\,.$ (59)
The shorthand notation $\{\bm{\xi}^{\gamma}\}$ stands for the average over all the replicas indexed by $\gamma$: $\{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}\}$, same is done for $\{\bm{\tau}^{\gamma}\}$ . Then the averaged partition function over the quenched disorder induced by the dataset is
$\left[Z^{n}\right]^{\bm{\mathcal{S}}}=\sum_{\bm{\mathcal{S}}}P_{\hat{\beta}}(\bm{\mathcal{S}})\ Z^{n}(\bm{\mathcal{S}})$ $=\sum_{\bm{\mathcal{S}}}\,\mathbb{E}_{\hat{\bm{\xi}}}\mathbb{E}_{\hat{\bm{\tau}}}\ z(\hat{\bm{\xi}})^{-M}\ \exp\left(\sqrt{\hat{\beta}/N}\sum_{ia}s_{i}^{a}\hat{\xi}_{i}\hat{\tau}_{a}\right)\ Z^{n}(\bm{\mathcal{S}})\,$, (60)
@@
[…]
$\hskip-34.14322ptz(\bm{\hat{\xi}})^{M}\approx\exp\left(-N\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}\hat{\beta}f_{T}(\bm{p}_{T}(\bm{\hat{\xi}}))\right)\,,\ \ z(\bm{\xi})^{M}\approx\exp\left(-N\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}\beta f_{S}(\bm{p}_{S}(\bm{\xi}))\right)$ (62)
@@
$z(\hat{\bm{\xi}})^{M}=\int\prod_{a}\left[\mathcal{D}{\bm{g}}^{a}\right]\,\mathbb{E}_{\bm{\hat{\tau}}}\exp$ $\left(\sqrt{\frac{\hat{\beta}}{N}}\sum_{i=1}^{N}\sum_{a=1}^{M}\sqrt{\Omega_{s}}\,g_{s,i}^{a}\hat{\tau}^{a}\hat{\xi}_{i}\right)\times$
$\times\prod_{i,a}2\cosh\left\{\sqrt{\frac{\hat{\beta}}{N}}\sqrt{\delta_{s}}\hat{\tau}^{a}\hat{\xi}_{i}\right\}$ (63)
$\hskip-62.59596pt\approx\mathbb{E}_{\bm{\hat{\tau}}}\,\exp\left(\frac{\hat{\beta}}{2N}\sum_{i=1}^{N}\sum_{a=1}^{M}\left(\hat{\tau}_{a}\hat{\xi}_{i}\right)^{2}\right)\,$, (64)
@@
$\delta\left(d_{\hat{\tau}}-\frac{1}{M}\sum_{a}\hat{\tau}_{a}^{2}\right)$ $=$ $\frac{1}{2\pi}\int d\hat{d}_{\hat{\tau}}\,\exp\left(-M\hat{d}_{\hat{\tau}}\left(d_{\hat{\tau}}-\frac{1}{M}\sum_{a}\hat{\tau}_{a}^{2}\right)\right)\,$, (65)
@@
$z(\hat{\bm{\xi}})^{M}$ $=$ $\int dd_{\hat{\tau}}\;\mathcal{D}(\hat{d}_{\hat{\tau}},d_{\hat{\tau}})\exp\left(\frac{\alpha\hat{\beta}}{2}d_{\hat{\tau}}\sum_{i}\hat{\xi}_{i}^{2}\right)$ (67)
$\approx$ $\exp\left(-N\hat{\beta}\operatorname{Extr}_{\bm{p}_{T}}f(\bm{p}_{T})\right)\,$, (68)
@@
$-\hat{\beta}f(\bm{p}_{T})=\alpha\left(-d_{\hat{\tau}}\hat{d}_{\hat{\tau}}+\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\hat{d}_{\hat{\tau}}+\log\mathbb{E}_{\hat{\tau}}\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\right)$ (69)
@@
$\hat{d}_{\hat{\tau}}=\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\,,\qquad\qquad d_{\hat{\tau}}=\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}=\frac{\mathbb{E}_{\hat{\tau}}\hat{\tau}^{2}\exp\left(\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\hat{\tau}^{2}\right)}{\mathbb{E}_{\hat{\tau}}\exp\left(\frac{\hat{\beta}}{2}d_{\hat{\xi}}(\hat{\bm{\xi}})\hat{\tau}^{2}\right)}\,.$ (70)
@@
$\hskip-42.67912pt\left[Z^{n}\right]^{\bm{\mathcal{S}}}$ $=$ $\sum_{\bm{\mathcal{S}}}\,\mathbb{E}_{\hat{\bm{\xi}}}\mathbb{E}_{\hat{\bm{\tau}}}\exp\left(N\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f(\bm{p}_{T})\right)\ \exp\left(\sqrt{\hat{\beta}/N}\sum_{ia}s_{i}^{a}\hat{\xi}_{i}\hat{\tau}_{a}\right)Z^{n}(\bm{\mathcal{S}})$ (71)
@@
The same problem of the teacher normalization term affects Eq.(58), due to the presence of $z(\{\bm{\xi}^{\gamma}\})$. As already stated in Eqs.(62) we want to write it in a variational principle manner. The steps to do that are the same of the one used for (63-70). We start computing the sum over the examples using their interpolated form, then we impose the self overlap with the Dirac delta
$\hskip-19.91684ptz(\{\bm{\xi}^{\gamma}\})^{M}=\int\prod_{a,\gamma}\left[\mathcal{D}\bm{g}^{a\gamma}\right]\,\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\exp$ $\left(\sqrt{\frac{\beta}{N}}\sum_{\gamma,i,a}\sqrt{\Omega_{s}}\,g_{s,i}^{a\gamma}{\tau}_{a}^{\gamma}\xi_{i}^{\gamma}\right)\times$
$\times\prod_{i,a,\gamma}2\cosh\left\{\sqrt{\frac{\beta}{N}}\sqrt{\delta_{s}}\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right\}$ (72)
$\approx\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\,\exp\left(\frac{\beta}{2N}\sum_{\gamma=1}^{n}\sum_{i=1}^{N}\sum_{a=1}^{M}\left(\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right)^{2}\right)\,$, (73)
$\approx\int\prod_{\gamma}\,dd_{\tau}^{\gamma}\,\mathcal{D}(\bm{\hat{d}_{\tau}},\bm{d_{\tau}})\exp\left(\frac{\alpha\beta}{2}\sum_{\gamma}d_{\tau}^{\gamma}d^{\gamma}({\bm{\xi}})\right)$ (74)
$\approx\exp\left(-N\beta\operatorname{Extr}_{\bm{p}_{S}}f(\bm{p}_{S})\right)\,$, (75)
@@
$\mathcal{D}(\bm{\hat{d}_{\tau}},\bm{d_{\tau}})$ $=$ $\sum_{\{\bm{\tau}^{\gamma}\}}\delta\left(\sum_{\gamma}d_{{\tau}}^{\gamma}-\frac{1}{M}\sum_{a,\gamma}(\tau_{a}^{\gamma})^{2}\right)$
$\propto$ $\sum_{\{\bm{\tau}^{\gamma}\}}\int\prod_{\gamma}d\hat{d}_{\tau}^{\gamma}\,\exp\left(-M\sum_{\gamma}\hat{d}_{\tau}^{\gamma}\left(d_{\tau}^{\gamma}-\frac{1}{M}\sum_{a}(\tau_{a}^{\gamma})^{2}\right)\right)\,$,
@@
$Z^{n}(\bm{\mathcal{S}})=\mathbb{E}_{\{\bm{\xi}^{\gamma}\}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\exp$ $\Bigg(M\sqrt{\frac{\beta}{N}}$ $\sum_{\gamma=1}^{n}p_{\tau}^{\gamma}(\tau)\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}p^{\gamma}(\xi)\tau^{\gamma}_{(1)}+$
$+$ $\sqrt{\frac{\beta}{N}}\sum_{\gamma=1}^{n}\sum_{i=2}^{N}\sum_{a=2}^{M}{s}_{i}^{a}\tau^{a,\gamma}\xi_{i}^{\gamma}+N\beta\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}f_{S}(\bm{p}_{S})\Bigg)\,.$
@@
$\hskip-19.91684pt-\beta f_{S}(\bm{p}_{S})=-\alpha\sum_{\gamma}d_{{\tau}}^{\gamma}\hat{d}_{{\tau}}^{\gamma}+\frac{\alpha\beta}{2}\sum_{\gamma}d^{\gamma}({\bm{\xi}}){d}_{{\tau}}^{\gamma}+\frac{1}{N}\log\mathbb{E}_{\{\tau^{\gamma}_{a}\}}\exp\left(\sum_{\gamma,a}\hat{d}_{{\tau}}^{\gamma}\sum_{a}(\tau^{\gamma}_{a})^{2}\right)$ (78)
$\hat{d}_{{\tau}}^{\gamma}=\frac{\beta}{2}d^{\gamma}({\bm{\xi}})\,,\qquad d^{\gamma}_{\hat{\tau}}=\langle{(\tau^{\gamma})}^{2}\rangle_{\tau}=\frac{\mathbb{E}_{\tau^{\gamma}}(\tau^{\gamma})^{2}\,\exp\left(\frac{\beta}{2}d^{\gamma}(\bm{\xi})(\tau^{\gamma})^{2}\right)}{\mathbb{E}_{\tau^{\gamma}}\exp\left(\frac{\beta}{2}d^{\gamma}(\bm{\xi})(\tau^{\gamma})^{2}\right)}\,.$ (79)
RS free energy and saddle point equations
@@
$\left[Z^{n}\right]^{\bm{\mathcal{S}}}=$ $\sum_{\{\xi_{(1)}^{\gamma}\}}\mathbb{E}_{\{\tau_{(1)}^{\gamma}\}}\sum_{\bm{s},\,\tilde{\bm{s}}}\exp{\left(M\sqrt{\frac{\beta}{N}}p_{\tau}^{\gamma}(\tau)\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}p(\xi)^{\gamma}\tau^{\gamma}_{(1)}\right)}\times$ (80)
$\times$ $\mathbb{E}_{\hat{\bm{\xi}}}\mathbb{E}_{\hat{\bm{\tau}}}\sum_{\bm{\mathcal{S}}}\exp\left(N\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f_{T}(\bm{p}_{T})+\sqrt{\frac{\hat{\beta}}{N}}\sum_{i=2}^{N}\sum_{a=2}^{M}{s}_{i}^{a}\hat{\tau}^{a}\hat{\xi}_{i}\right)\times$ (81)
$\times$ $\sum_{\{\bm{\xi}^{\gamma}\}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\;\exp\left(N\beta\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}f_{S}(\bm{p}_{S})+\sqrt{\frac{\beta}{N}}\sum_{i=2}^{N}\sum_{a=2}^{M}\sum_{\gamma=1}^{n}{s}_{i}^{a}\tau^{a,\gamma}\xi_{i}^{\gamma}\right)\;.$ (82)
@@
$\int\,\prod_{\gamma}dp^{\gamma}\,\prod_{\gamma}dp_{\tau}^{\gamma}\,\mathcal{D}(\bm{p},\bm{p}_{\tau})\sum_{\{\xi_{(1)}^{\gamma}\}}\mathbb{E}_{\{\tau_{(1)}^{\gamma}\}}\exp{\left(M\sqrt{\frac{\beta}{N}}\sum_{\gamma}p_{\tau}^{\gamma}\xi_{(1)}^{\gamma}+N\sqrt{\frac{\beta}{N}}\sum_{\gamma}p^{\gamma}\tau^{\gamma}_{(1)}\right)}$
$=$ $\int\,\prod_{\gamma}dp^{\gamma}\,\prod_{\gamma}dp_{\tau}^{\gamma}\,\mathcal{D}(\bm{p},\bm{p}_{\tau})\exp{\left(\frac{N\beta}{2}\alpha^{2}\Omega\sum_{\gamma}(p_{\tau}^{\gamma})^{2}+\frac{N\beta}{2}\Omega_{\tau}\sum_{\gamma}(p^{\gamma})^{2}\right)}\,$, (83)
@@
$\propto\sum_{\bm{s},\,\tilde{\bm{s}}}\int\,\prod_{\gamma}d\hat{p}^{\gamma}\prod_{\gamma}d\hat{p}_{\tau}^{\gamma}\exp\left(-N\sum_{\gamma}\hat{p}_{\gamma}\left(p_{\gamma}-\frac{1}{N}\sum_{i}s_{i}\xi^{\gamma}_{i}\right)\right)\times$
$\hskip 105.2751pt\times\exp\left(-M\sum_{\gamma}\hat{p}^{\gamma}_{\tau}\left(p^{\gamma}_{\tau}-\frac{1}{M}\sum_{a}\tilde{s}_{a}\tau^{\gamma}_{a}\right)\right)\;.$